# CoPL: Collaborative Preference Learning for Personalizing LLMs

### v1

@@ Abstract
Personalizing large language models (LLMs) is important for aligning outputs with diverse user preferences, yet existing methods struggle with flexibility and generalization. We propose CoPL (Collaborative Preference Learning), a graph-based collaborative filtering framework that models user-response relationships to enhance preference estimation, particularly in sparse annotation settings. By integrating a mixture of LoRA experts, CoPL efficiently fine-tunes LLMs while dynamically balancing shared and user-specific preferences. Additionally, an optimization-free adaptation strategy enables generalization to unseen users without fine-tuning. Experiments on UltraFeedback-P demonstrate that CoPL outperforms existing personalized reward models, effectively capturing both common and controversial preferences, making it a scalable solution for personalized LLM alignment.
@@
Youngbin Choi1, Seunghyuk Cho1, Minjong Lee 2, MoonJeong Park1, Yesong Ko2, Jungseul Ok1,2, Dongwoo Kim1,2 1Graduate School of Artificial Intelligence, POSTECH, 2Department of Computer Science and Engineering, POSTECH, {choi.youngbin, shhj1998, minjong.lee, mjeongp, yesong.ko, jungseul, dongwoo.kim}@postech.ac.kr
Correspondence to: Dongwoo Kim ¡dongwoo.kim@postech.ac.kr¿
@@ 1 Introduction
Large language models (LLMs) have rapidly expanded across diverse applications, from customer service and tutoring to creative content generation Shi et al. (2024); Molina et al. (2024); Venkatraman et al. (2024). As increasing numbers of users with varied backgrounds interact with LLMs, accounting for diverse preferences has become essential. Most reward models rely on the Bradley-Terry-Luce (BTL) framework (Bradley and Terry, 1952), which learns preferences from pairwise comparisons provided by human annotators. However, earlier studies largely assumed a single, uniform preference and neglected the diversity of user preferences (Siththaranjan et al., 2024; Li et al., 2024). This limitation has led to growing interest in personalized reward models (Sorensen et al., 2024).
@@
There are two different approaches to utilizing the BTL framework for personalized reward models. The first approach has explored combining multiple reward models, each trained for a specific preference and later aggregated (Jang et al., 2023; Oh et al., 2024). However, this approach relies on pre-trained models for different preference types, reducing flexibility. Another line of work introduces user-specific latent variables into a single BTL framework, learning personalized representations from user annotations (Chen et al., 2024a; Poddar et al., 2024; Li et al., 2024). While this method captures individual preferences, the latent variable model does not explicitly account for relationships between users sharing similar responses. As a result, it struggles to generalize in sparse annotation settings.
@@
To address these limitations, we propose Collaborative Preference Learning (CoPL), which constructs a user-response bipartite preference graph from pairwise annotations and uses a graph-based collaborative filtering (GCF) framework for personalized reward modeling. Unlike approaches that model each user separately, GCF on the graph structure allows preference signals to propagate across users, enabling to exploit multi-hop relationships among users and responses (Wang et al., 2019; He et al., 2020). As a result, CoPL can capture diverse preferences of users even in sparse annotation settings.
@@
Based on the user embedding, we develop an LLM-based reward model that can predict the preference score of a user given input text. We adopt the mixture of LoRA experts (MoLE) (Chen et al., 2023, 2024c; Liu et al., 2024) that allows parameter efficient fine-tuning while routing different users to different paths based on the learned embedding. Specifically, we develop a user preference-aware gating function that dynamically selects the experts in the forward pass, making the LLM predict a personalized preference.
@@ 2 Related Work
In this section, we summarize relevant lines of research, such as personalized alignment and preference learning with sparse interactions.
Personalized alignment.
Preference learning with sparse interactions.
@@ 3 Problem Formulation
Through the maximum likelihood estimation with preference data for all users, one can learn the reward function $f$ to make the reward function align with user preference. In the case of the universal preference model, user $u$ is ignored in Eq. 1 (Chen et al., 2024b; Achiam et al., 2023; Dai et al., 2023; Bai et al., 2022). In practice, the user $u$ is replaced by a user embedding (Poddar et al., 2024; Li et al., 2024; Chen et al., 2024a).
@@ 4 Method
In this section, we describe our Collaborative Preference Learning (CoPL). We first learn user embeddings based on GCF with the preference data. We then train the reward model based on the learned user embeddings. Finally, we provide an optimization-free adaptation strategy to obtain embeddings of users who are unseen during training.
@@ 4.1 User Representation Learning
The preference dataset for all users can be naturally converted into a bipartite graph, where each user and response is represented as a node, and an edge between a user and a response represents the user’s preference over the response. The edge can have two different types: positive or negative, indicating whether a user prefers the response or not.
@@
$e_{u}^{(\ell+1)}={\psi}\bigl{(}m^{(\ell)}_{u}\bigr{)},$ (3)
where $\psi(\cdot)$ is a non-linear activation. The response embedding $e_{r}^{(\ell)}$ is updated with analogous process. We randomly initialize the user and response embeddings at the first layer and then fine-tune the embeddings through training. The update steps for the response embeddings are provided in Appendix A.
@@ 4.3 Optimization-free User Adaptation
While we can predict a preference score of unseen responses for a known user, the reward model trained in Section 4.2 cannot be used to predict the preference of users who have not been observed during training. To estimate the embeddings of unseen users, we propose an optimization-free adaptation approach.
@@
[…]
@@ 5 Experiments
In this section, we aim to show whether reward models can accurately learn user preferences in sparse annotation scenarios. Specifically, we examine situations where many users contribute only a few annotated pairs.
@@ 5.1 Experimental Settings
We employ the UltraFeedback-P (UF-P) dataset (Poddar et al., 2024), which is explicitly designed to capture diverse user preferences from UltraFeedback (Cui et al., 2023). Unlike traditional reward modeling datasets that assume a single dominant preference, UF-P explicitly builds diverse preference groups through fine-grained scores across multiple preference attributes about response from UltraFeedback.
UF-P is created by grouping users based on distinct preference priorities, including helpfulness, honesty, instruction-following, and truthfulness. This dataset consists of two environments, each with a different number of groups. First, UF-P-2 consists of two user groups, each prioritizing either helpfulness or honesty. UF-P-4 expands to four groups, each concentrating on a different attribute. We provide a detailed explanation of the construction of the UF-P dataset from UltraFeedback in Section C.1.
While UF-P supports personalized reward modeling, it does not inherently reflect scenarios where a large number of users each provides only a handful of annotations. To reflect our target scenario, we generate a modified version of UF-P with 10,000 users evenly distributed across different preference groups and a survey set of 25,993 pairs.
Specifically, we construct four experimental environments based on UF-P-2 and UF-P-4:
  • UF-P-2-ALL: In two preference groups, each user contributes exactly 8 annotations.
  • UF-P-2-AVG: In two preference groups, each user contributes 8 annotations on average.
  • UF-P-4-ALL: In four preference groups, each user contributes exactly 16 annotations.
  • UF-P-4-AVG: In four preference groups, each user contributes 16 annotations on average.
For UF-P-2-AVG and UF-P-4-AVG, we randomly sample the number of annotations from a uniform distribution over $1\sim 15$ and $1\sim 31$, respectively.
Gemma-2b-it Gemma-7b-it
UF-P-2 UF-P-4 UF-P-2 UF-P-4
@@
Seen Oracle $64.53_{\pm 0.14}$ $64.53_{\pm 0.14}$ $61.52_{\pm 0.13}$ $61.52_{\pm 0.13}$ $66.80_{\pm 0.17}$ $66.80_{\pm 0.17}$ $62.17_{\pm 0.09}$ $62.17_{\pm 0.09}$
Uniform $61.82_{\pm 0.16}$ $61.82_{\pm 0.16}$ $56.15_{\pm 0.22}$ $56.15_{\pm 0.22}$ $61.96_{\pm 0.07}$ $61.96_{\pm 0.07}$ $56.80_{\pm 0.12}$ $56.80_{\pm 0.12}$
I2E $61.48_{\pm 0.18}$ $61.49_{\pm 0.70}$ $57.21_{\pm 0.37}$ $57.44_{\pm 0.37}$ $62.10_{\pm 0.28}$ $61.43_{\pm 0.23}$ $57.90_{\pm 0.21}$ $58.50_{\pm 0.09}$
@@
VPL $61.11_{\pm 0.16}$ $61.86_{\pm 0.84}$ $56.04_{\pm 1.71}$ $56.77_{\pm 0.38}$ $62.39_{\pm 0.10}$ $62.59_{\pm 0.24}$ $58.87_{\pm 0.25}$ $57.55_{\pm 1.00}$
PAL $59.95_{\pm 0.04}$ $61.53_{\pm 0.22}$ $56.95_{\pm 0.13}$ $57.37_{\pm 0.14}$ $62.59_{\pm 0.06}$ $62.47_{\pm 0.13}$ $57.17_{\pm 0.22}$ $56.27_{\pm 0.13}$
CoPL $\textbf{63.81}_{\pm 0.16}$ $\textbf{63.45}_{\pm 0.38}$ $\textbf{62.57}_{\pm 0.38}$ $\textbf{62.08}_{\pm 0.27}$ $\textbf{63.90}_{\pm 0.07}$ $\textbf{63.48}_{\pm 0.13}$ $\textbf{62.90}_{\pm 0.05}$ $\textbf{61.93}_{\pm 0.02}$
Unseen Oracle $64.66_{\pm 1.10}$ $64.66_{\pm 1.10}$ $61.33_{\pm 0.35}$ $61.33_{\pm 0.35}$ $67.43_{\pm 0.65}$ $67.43_{\pm 0.65}$ $62.01_{\pm 0.04}$ $62.01_{\pm 0.04}$
Uniform $62.82_{\pm 0.59}$ $62.82_{\pm 0.59}$ $55.65_{\pm 0.61}$ $55.65_{\pm 0.61}$ $62.23_{\pm 0.06}$ $62.23_{\pm 0.06}$ $57.02_{\pm 0.27}$ $57.02_{\pm 0.27}$
I2E $61.67_{\pm 0.82}$ $59.52_{\pm 0.51}$ $56.42_{\pm 0.41}$ $56.75_{\pm 0.68}$ $62.62_{\pm 0.95}$ $61.88_{\pm 0.21}$ $57.62_{\pm 0.92}$ $58.12_{\pm 0.98}$
@@
VPL $60.83_{\pm 0.40}$ $62.62_{\pm 0.49}$ $54.03_{\pm 1.54}$ $56.13_{\pm 0.57}$ $62.69_{\pm 0.99}$ $63.67_{\pm 0.12}$ $58.49_{\pm 1.22}$ $56.85_{\pm 0.84}$
PAL $59.83_{\pm 0.69}$ $61.71_{\pm 0.31}$ $57.07_{\pm 0.22}$ $57.13_{\pm 0.33}$ $63.08_{\pm 0.73}$ $62.52_{\pm 0.58}$ $57.15_{\pm 0.48}$ $56.44_{\pm 0.67}$
CoPL $\textbf{63.92}_{\pm 0.54}$ $\textbf{63.26}_{\pm 0.51}$ $\textbf{61.62}_{\pm 0.10}$ $\textbf{61.97}_{\pm 0.35}$ $\textbf{64.08}_{\pm 0.71}$ $\textbf{64.38}_{\pm 1.00}$ $\textbf{62.77}_{\pm 1.32}$ $\textbf{62.08}_{\pm 0.64}$
Table 1: Accuracy of reward models on unseen annotated pairs. The Seen user results report performance for all users encountered during training in the upper block of the table. The Unseen user results report performance for 100 new users, evenly distributed across preference groups. Unseen users provide 8 annotations under UF-P-2-ALL/AVG and 16 annotations under UF-P-4-ALL/AVG. Bold represents the best result, except with Oracle. All experiments run on three seeds.
(a) I2E
[…]
(c) VPL
(d) PAL
(e) CoPL
Figure 1: T-SNE visualization of seen user embeddings in UF-P-4-AVG with gemma-2b-it. Points are colored by their preference group. Our method clusters users in the same group more effectively, whereas other baselines fail to cluster users by their preference groups in user embedding space.
Since UF-P-4 encompasses a broader range of preferences, users provide more annotations to capture this added complexity. These configurations enable us to rigorously evaluate how reward models perform under sparse user annotations, a critical challenge for large-scale personalized alignment in practical settings.
Notably, our experimental environments remain consistent with previous work (Poddar et al., 2024), but more closely mirror our target environments. Specifically, Poddar et al. (2024) infers user preferences from a small, predefined pool of unannotated pairs, so all users must be evaluated within that limited query set. In contrast, we consider a much broader range of unannotated pairs, allowing the model to capture preferences across diverse contexts and better adapt to real-world personalized alignment scenarios.
@@
Table 2: Accuracy of reward models on UF-P-2-ALL with gemma-2b-it, broken down by pair type. Common refers to pairs for which the two preference groups provide the same preference label, Controversial refers to pairs labeled differently by the two groups, and Total encompasses all pairs. These categories reflect how diverse user preferences affect the performance of reward models. Bold represents the best result, except with Oracle.
Training and evaluation details.
For reward function training, we utilize two LLM backbones: gemma-2b-it and gemma-7b-it (Team et al., 2024). Our model uses one shared LoRA, eight LoRA experts, each with a rank of eight, and a two-layer MLP for the gating function. The other baselines, e.g., Uniform, I2E, VPL, PAL, and Oracle, use a LoRA rank of 64. Other training details, such as hyper-parameters and model architecture, are provided in Section C.2.
We report reward model accuracy on unseen test pairs that are not in the survey set. We define a correct prediction as assigning a higher score to the preferred response. We evaluate performance for both seen and unseen users. For seen user experiments, each user is assigned 10 test pairs, and accuracy is calculated over all seen users. We fix the number of unseen users at 100, evenly distributed across preference groups. To adapt the reward model for each unseen user, we provide 8 annotations in UF-P-2-ALL/AVG and 16 annotations in UF-P-4-ALL/AVG, followed by evaluation on 50 test pairs per unseen user. CoPL uses 2-hop neighbors for unseen user adaptation.
(a) UF-P-2-AVG
(b) UF-P-4-AVG
Figure 2: Accuracy of unseen user adaptation as the number of provided annotation sets increases, evaluated on UF-P-2/4-AVG with gemma-2b-it. 2-hop and 4-hop indicates 2-hop and 4-hop adaptation, respectively.

### v2

@@ Abstract
Personalizing large language models (LLMs) is important for aligning outputs with diverse user preferences, yet existing methods struggle with flexibility and generalization. We propose CoPL (Collaborative Preference Learning), a graph-based collaborative filtering framework that models user-response relationships to enhance preference estimation, particularly in sparse annotation settings. By integrating a mixture of LoRA experts, CoPL efficiently fine-tunes LLMs while dynamically balancing shared and user-specific preferences. Additionally, an optimization-free adaptation strategy enables generalization to unseen users without fine-tuning. Experiments on TL;DR, UltraFeedback-P, and PersonalLLM datasets demonstrate that CoPL outperforms existing personalized reward models, effectively capturing both common and controversial preferences, making it a scalable solution for personalized LLM alignment. The code is available at https://github.com/ml-postech/CoPL.
@@
Youngbin Choi1, Seunghyuk Cho1, Minjong Lee 2, MoonJeong Park1, Yesong Ko2, Jungseul Ok1,2, Dongwoo Kim1,2,* 1Graduate School of Artificial Intelligence, POSTECH, 2Department of Computer Science and Engineering, POSTECH, {choi.youngbin, shhj1998, minjong.lee, mjeongp, yesong.ko, jungseul, dongwoo.kim}@postech.ac.kr
footnotetext: Correspondence to: Dongwoo Kim <dongwoo.kim@postech.ac.kr>
@@ 1 Introduction
Large language models (LLMs) have rapidly expanded across diverse applications, from customer service and tutoring to creative content generation Shi et al. (2024); Molina et al. (2024); Venkatraman et al. (2024). As increasing numbers of users with varied backgrounds interact with LLMs, accounting for diverse preferences has become essential. Most reward models rely on the Bradley-Terry-Luce (BTL) framework (Bradley and Terry, 1952), which learns preferences from pairwise comparisons provided by human annotators. However, earlier studies largely assumed a single, uniform preference and neglected the diversity of user preferences (Siththaranjan et al., 2024; Li et al., 2024, 2025). This limitation has led to growing interest in personalized reward models (Sorensen et al., 2024; Liu et al., 2025; Guan et al., 2025).
(a) VPL
(b) CoPL
Figure 1: T-SNE visualization of seen user embeddings in UF-P-4 (AVG) with gemma-2b-it. Points are colored by their preference group. Our method clusters users in the same group more effectively. T-SNE visualizations of other baselines are provided in Fig.˜A1.
@@
There are two different approaches to utilizing the BTL framework for personalized reward models. The first approach has explored combining multiple reward models, each trained for a specific preference and later aggregated (Jang et al., 2023; Oh et al., 2024). However, this approach relies on pre-trained models for different preference types, reducing flexibility. Another line of work introduces user-specific latent variables into a single BTL framework, learning personalized representations from user annotations (Chen et al., 2024a; Poddar et al., 2024; Li et al., 2024; Barreto et al., 2025). While this method captures individual preferences, the latent variable model does not explicitly account for relationships between users sharing similar responses. As a result, it struggles to generalize in sparse annotation settings.
@@
To address these limitations, we propose Collaborative Preference Learning (CoPL), which constructs a user-response bipartite preference graph from pairwise annotations and uses a graph-based collaborative filtering (GCF) framework for personalized reward modeling. Unlike approaches that model each user separately, GCF on the graph structure allows preference signals to propagate across users, enabling to exploit multi-hop relationships among users and responses (Wang et al., 2019; He et al., 2020). CoPL can capture diverse preferences of users even in sparse annotation settings.
When annotations are sparse, latent-variable methods face significant challenges, as the scarcity of supervisory signals makes it difficult for randomly initialized user representation encoders to converge toward semantically meaningful representations. As a result, users with similar underlying preferences can sometimes be mapped to distant points in the latent space if their annotated response pair sets do not overlap. In such cases, sparse supervision may cause semantically similar users to appear unrelated in the learned embedding space. For instance, consider three users: user 1 annotates the pairs (a,b),(c,d){(a,b),(c,d)}, user 2 annotates (c,d),(e,f){(c,d),(e,f)}, and user 3 annotates (e,f),(g,h){(e,f),(g,h)} with the same preference. Although user 1 and user 3 exhibit similar preferences, the lack of overlapping annotations provides no direct signal for aligning their representations. CoPL addresses this issue by constructing a user–response bipartite graph and propagating preference signals through multi-hop message passing. This mechanism enables the alignment of users with disjoint annotation sets, such as user 1 and user 3, thereby providing better data efficiency and generalization. Fig.˜1 illustrates that, under sparse annotation, CoPL produces embedding spaces in which users with identical preferences are more coherently aligned.
@@
Based on the user embedding, we develop an LLM-based reward model that can predict the preference score of a user given input text. We adopt the mixture of LoRA experts (MoLE) (Chen et al., 2023, 2024c; Liu et al., 2024) that allows parameter-efficient fine-tuning while routing different users to different paths based on the learned embedding. Specifically, we develop a user preference-aware gating function that dynamically selects the experts in the forward pass, making the LLM predict a personalized preference.
@@ 2 Related Work
Alignment has emerged as a crucial strategy for mitigating undesirable outcomes (Dai et al., 2023; Yang et al., 2024a). Previous research has often focused on the average preference of annotators (Achiam et al., 2023), ignoring the diverse preferences. To address preference diversity, recent works (Jang et al., 2023; Oh et al., 2024; Yang et al., 2024b) view this problem as a soft clustering problem, where user-specific preferences are treated as mixtures of predefined preference types. Although this approach effectively handles diverse preferences, it relies on specifying several preference types in advance.
Another line of work introduces a user latent variable in the BTL framework (Poddar et al., 2024; Li et al., 2024; Chen et al., 2024a). The main challenge lies in obtaining user representations. One approach is to treat each user embedding as learnable parameters (Li et al., 2024; Chen et al., 2024a), and the other strategy is to train an encoder that infers embeddings from the set of annotated pairs provided by each user (Poddar et al., 2024).
We also discuss preference learning with sparse interactions, closely related to our approach, in Appendix˜C.
@@ 3 Problem Formulation
Through the maximum likelihood estimation with preference data for all users, one can learn the reward function $f$ to make the reward function align with user preference. In the case of the universal preference model, user $u$ is ignored in Eq.˜1 (Chen et al., 2024b; Achiam et al., 2023; Dai et al., 2023; Bai et al., 2022). In practice, the user $u$ is replaced by a user embedding (Poddar et al., 2024; Li et al., 2024; Chen et al., 2024a).
@@ 4 Method
Figure 2: An overview of CoPL. To learn user representations, the GCF model is trained on a user-response bipartite graph. To build a personalized reward model, CoPL uses the learned representations to select a user-specific expert from MoLE, enabling effective modeling of diverse preferences.
In this section, we describe our Collaborative Preference Learning (CoPL). Our approach consists of three steps: learning user representations given preference data, construction of personalized reward models, and adaptation to unseen (new) users at test time. Figure˜2 illustrates the first two steps, and Figure 3 the last step.
@@ 4.1 User Representation Learning
The preference dataset for all users can be naturally converted into a bipartite graph, where each user and response is represented as a node, and an edge between a user and a response represents the user’s preference over the response, as illustrated in Fig.˜2. The edge can have two different types: positive or negative, indicating whether a user prefers the response or not.
@@
[…]
where $\psi(\cdot)$ is a non-linear activation. The response embedding $e_{r}^{(\ell)}$ is updated with analogous process. We randomly initialize the user and response embeddings at the first layer and then fine-tune the embeddings through training. The update steps for the response embeddings are provided in Appendix˜A.
@@ 4.2 Personalized Reward Model with User Representations
Figure 3: Illustration of unseen user adaptation. Blue nodes are users who have similar preferences to $u^{*}$, and red nodes are users who have dissimilar preferences.
@@ 4.3 Optimization-free User Adaptation
While we can predict a preference score of unseen responses for a known user, the reward model trained in Section˜4.2 cannot be used to predict the preference of users who have not been observed during training. To estimate the embeddings of unseen users, we propose an optimization-free adaptation approach.
@@
where $\mathcal{N}^{+}_{u{{}^{*}}}(k)$ is a set of $k$-hop neighborhood111kk must be an even number to aggregate only the user embeddings. of user $u^{*}$ connected by only positive edges, and $w_{u,u^{*}}$ is a normalized alignment score between $u$ and $u^{*}$. The normalized alignment score $w_{u,u^{*}}$ is defined as
@@ 5 Experiments
In this section, we empirically verify the performance of CoPL across various scenarios.
Dataset TL;DR UF-P-2 UF-P-4 PersonalLLM
Size of survey set 19,824 25,993 25,993 14,435
# of preference groups 2 2 4 $\infty$
# of annotations per user 8 8 16 16
# of users per group 5,000 5,000 2,500 -
Table 1: Statistics of the datasets. We report the average number of annotations per user. All users have different preferences in PersonalLLM.
@@ 5.1 Experimental Settings
We employ three datasets, including TL;DR (Stiennon et al., 2020; Chen et al., 2024a), UltraFeedback-P (UF-P) (Poddar et al., 2024), and PersonalLLM (Zollo et al., 2024), that explicitly capture diverse user preferences rather than assuming a single dominant preference. We briefly describe the key characteristics of these datasets below.
Following prior work (Chen et al., 2024a; Li et al., 2024), we define two user groups in the TL;DR dataset: one group prefers short summaries, and the other favors long summaries. We create two environments with the UF-P dataset: UF-P-2, dividing users into two groups based on their preference, and UF-P-4, dividing users into four groups. In PersonalLLM (Zollo et al., 2024), user preferences are modeled as a mixture of four preference dimensions where weight vectors are drawn from a Dirichlet distribution with $\alpha=0.1$. Additional details on their construction and properties can be found in Section˜D.1.
TL;DR UF-P-2 UF-P-4 PersonalLLM
@@
Seen
G-Oracle $73.06_{\pm 0.23}$ $73.06_{\pm 0.23}$ $64.53_{\pm 0.14}$ $64.53_{\pm 0.14}$ $61.52_{\pm 0.13}$ $61.52_{\pm 0.13}$ N/A N/A
Uniform $49.62_{\pm 0.09}$ $49.62_{\pm 0.09}$ $61.82_{\pm 0.16}$ $61.82_{\pm 0.16}$ $56.15_{\pm 0.22}$ $56.15_{\pm 0.22}$ $62.91_{\pm 0.07}$ $62.91_{\pm 0.07}$
I2E $49.93_{\pm 0.23}$ $49.74_{\pm 0.06}$ $61.48_{\pm 0.18}$ $61.49_{\pm 0.70}$ $57.21_{\pm 0.37}$ $57.44_{\pm 0.37}$ $65.74_{\pm 0.04}$ $65.77_{\pm 0.05}$
@@
VPL $49.52_{\pm 0.14}$ $49.44_{\pm 0.21}$ $61.11_{\pm 0.16}$ $61.86_{\pm 0.84}$ $56.04_{\pm 1.71}$ $56.77_{\pm 0.38}$ $70.84_{\pm 0.18}$ $67.95_{\pm 0.21}$
PAL $2_{\pm 0.13} 50.15±0.1$$5_{\pm 0.15}$ $59.95_{\pm 0.04}$ $61.53_{\pm 0.22}$ $56.95_{\pm 0.13}$ $57.37_{\pm 0.14}$ $6.25_{\pm 0.35} 66$$6.29_{\pm 0.06}$
CoPL $\mathbf{96.58}_{\pm 0.09}$ $\mathbf{96.19}_{\pm 0.02}$ $\mathbf{63.81}_{\pm 0.16}$ $\mathbf{63.45}_{\pm 0.38}$ $\mathbf{62.57}_{\pm 0.38}$ $\mathbf{62.08}_{\pm 0.27}$ $\mathbf{74.85}_{\pm 0.17}$ $\mathbf{74.37}_{\pm 0.03}$
Unseen
G-Oracle $72.55_{\pm 1.79}$ $72.55_{\pm 1.79}$ $64.66_{\pm 1.10}$ $64.66_{\pm 1.10}$ $61.33_{\pm 0.35}$ $61.33_{\pm 0.35}$ N/A N/A
Uniform $50.11_{\pm 0.36}$ $50.11_{\pm 0.36}$ $62.82_{\pm 0.59}$ $62.82_{\pm 0.59}$ $5.65_{\pm 0.61} 55$$5.65_{\pm 0.61}$ $62.97_{\pm 0.07}$ $62.97_{\pm 0.07}$
I2E $49.85_{\pm 0.38}$ $49.16_{\pm 0.82}$ $61.67_{\pm 0.82}$ $59.52_{\pm 0.51}$ $56.42_{\pm 0.41}$ $56.75_{\pm 0.68}$ $5.79_{\pm 0.18} 66$$6.11_{\pm 0.24}$
@@
VPL $49.40_{\pm 0.88}$ $49.31_{\pm 0.57}$ $60.83_{\pm 0.40}$ $62.62_{\pm 0.49}$ $54.03_{\pm 1.54}$ $56.13_{\pm 0.57}$ $71.31_{\pm 0.58}$ $68.55_{\pm 0.47}$
PAL $49.48_{\pm 0.86}$ $49.64_{\pm 0.55}$ $59.83_{\pm 0.69}$ $61.71_{\pm 0.31}$ $57.07_{\pm 0.22}$ $57.13_{\pm 0.33}$ $5.94_{\pm 0.11} 66$$6.40_{\pm 0.03}$
CoPL $\mathbf{96.71}_{\pm 0.25}$ $\mathbf{96.21}_{\pm 0.14}$ $\mathbf{63.92}_{\pm 0.54}$ $\mathbf{63.26}_{\pm 0.51}$ $\mathbf{61.62}_{\pm 0.10}$ $\mathbf{61.97}_{\pm 0.35}$ $\mathbf{75.69}_{\pm 0.22}$ $\mathbf{75.49}_{\pm 0.03}$
Table 2: Accuracy of reward models on unseen annotated pairs. The results report performance on Seen users encountered during training and on Unseen users. Bold represents the best result, except for G-Oracle. These results are based on gemma-2b-it. Additional results using gemma-7b-it and gemma2-27b-it are represented in Table˜A1 and Table˜A3, respectively.
We divide $10,000$ users evenly into the predefined number of preference groups. For all datasets, we curate two different versions, denoted as ALL and AVG, representing two different annotation sampling strategies. For TL;DR/UF-P-2 (ALL), each user provides exactly 8 annotations, while for TL;DR/UF-P-2 (AVG), each user’s annotation count is uniformly sampled from 1 to 15, averaging to 8. Similarly, in UF-P-4/PersonalLLM (ALL), each user provides exactly 16 annotations, and in UF-P-4/PersonalLLM (AVG), the count is uniformly sampled from 1 to 31, averaging to 16. Table 1 summarizes the key statistics.
@@
Training and evaluation details.
For reward function training, we utilize two LLM backbones: gemma-2b-it and gemma-7b-it (Team et al., 2024a). Our model uses one shared LoRA, eight LoRA experts, each with a rank of eight, and a two-layer MLP for the gating function. The other baselines, e.g., Uniform, I2E, VPL, PAL, and G-Oracle, use a LoRA rank of 64. Other training details, such as hyper-parameters and model architecture, are provided in Section˜D.2. All experiments, including additional analysis, are repeated three times with different seeds.
We report reward model accuracy on unseen test pairs that are not in the survey set. We evaluate performance for both seen and unseen users. For seen user experiments, each user is assigned 10 test pairs, and accuracy is calculated over all seen users. We fix the number of unseen users at 100, evenly distributed across preference groups. To adapt the reward model for each unseen user, we provide 8 annotations in TL;DR/UF-P-2 (ALL/AVG) and 16 annotations in UF-P-4/PersonalLLM (ALL/AVG), followed by evaluation on 50 test pairs per unseen user. CoPL uses 2-hop neighbors for unseen user adaptation.
@@
Table 3: Accuracy of reward models on UF-P-2 (ALL) with gemma-2b-it, broken down by pair type. Common refers to pairs for which the two preference groups provide the same preference label, Controversial refers to pairs labeled differently by the two groups, and Total encompasses all pairs. These categories reflect how diverse user preferences affect the performance of reward models. Bold represents the best result, except with G-Oracle.